OpenStudy (anonymous):

Find The number of complex zeros that F(x) has in F(x)=x^7+x^9+x^12-2x^2

OpenStudy (anonymous):

apparently it has 8, but how you are supposed to do this i have no idea

OpenStudy (anonymous):

what method were you supposed to use?

OpenStudy (anonymous):

Any

OpenStudy (anonymous):

How did you find it will have 8?? @satellite73

OpenStudy (anonymous):

it will have 8 complex roots

OpenStudy (anonymous):

you have a twelfth degree polynomial for heaven sake we can factor out \(x^2\) but you still have a tenth degree polynomial how can you do it? i have no idea

OpenStudy (anonymous):

but i will be quiet and learn if @k.rajabhishek has a way to know it without graphing

OpenStudy (anonymous):

since the highest degree is 12 thus it will have 12 zeros

OpenStudy (anonymous):

And out of 12 zeros how many are complex??? @k.rajabhishek

OpenStudy (anonymous):

got that and 0 is a zero with multiplicity 2

OpenStudy (anonymous):

by using descarte's rule we will have 2 negative and 1 positive zeros also 0 is one of the zero of equation

OpenStudy (anonymous):

thus a total of 4 real zeros and then subtracting 4 by 12 we have 8 complex roots

OpenStudy (anonymous):

wow that is impressive

OpenStudy (anonymous):

except i think that there is only one negative zero 0 is a zero with multiplicity 2 descartes rule of sign is going to tell us there is one positive root right?

OpenStudy (anonymous):

there is some mistake with this method

OpenStudy (anonymous):

ok descartes rule of sign will give it, but we have to change it a bit

OpenStudy (anonymous):

first we need to write in standard form then we need to get rid of the \(x^2\) and then descartes rule of sign will tell use there is one positive root and one negative root, leaving 8 complex roots

OpenStudy (anonymous):

@Theanitrix you got this? we can walk through it if you like

OpenStudy (anonymous):

\[F(x)=x^7+x^9+x^{12}-2x^2\] \[F(x)=x^2(x^{10}+x^7+x^5-2)\] now we know it has a zero at \(x=0\) with multiplicity 2 that leaves 10 more to find

OpenStudy (anonymous):

start with\(f(x)=x^{10}+x^7+x^5-2\) there is one change in sign, from \(1x^5\) to \(-2\) telling us that there is one positive root

OpenStudy (anonymous):

then \(f(-x)=x^{10}-x^7-x^5-2\) and this has one change in sign from \(1x^{10}\) to \(-x^7\) telling us there is one negative root

OpenStudy (anonymous):

that leaves 8 complex roots @k.rajabhishek had method but a mistake in the execution

OpenStudy (anonymous):

@satellite73 I am not getting it.. Will you clear my doubt?? How to find the positive and negative roots according to the sign change as you did..??

OpenStudy (matheducatormcg):

fantastic discussion on complex roots of polynomials!

OpenStudy (anonymous):

there is one change in sign of the coefficients that means there is one positive root, if there were two then there could be two or none, you count down by twos but you can't count down by twos if there is only one similarly there is one change in sign of the coefficients of \(f(-x\) so again there must be one negative root ten in total, leaves 8 complex roots descartes rule of sign explained well here http://www.purplemath.com/modules/drofsign.htm